The actual selling expenses incurred in March 2020 by Fallon Company are as follows.

Fixed Expenses

Sales salaries$34,900

Depreciation 6,600

Insurance1,300

Variable Expenses

Sales commissions$11,340

Advertising 12,280

Travel 6,960

Delivery 3,600


(a) Prepare a flexible budget performance report for March, assuming that March sales were $174,000. Variable costs and their percentage relationship to sales are sales commissions 6%, advertising 7%, traveling 4%, and delivery 2%. Fixed selling expenses will consist of sales salaries $34,900, Depreciation on delivery equipment $6,600, and insurance on delivery equipment $1,300. (List variable costs before fixed costs.)

(b) Prepare a flexible budget performance report, assuming that March sales were $181,900. (List variable costs before fixed costs.)
Answer #1

a)

Fallon Company
Selling Expense Flexible Budget Report
For the Month Ended March 31, 2020
Budget Actual Difference
Sales $174000 $174000 0 NONE
Less: Variable costs:
Sales commissions (174000*6%)= 10440 11340 (10440-11340)= 900 U
Advertising (174000*7%)= 12180 12280 (12180-12280)= 100 U
Travel (174000*4%)= 6960 6960 0 NONE
Delivery (174000*2%)= 3480 3600 (3480-3600)= 120 U
Total variable costs $33060 $34180 (33060-34180)= 1120 U
Less: Fixed costs
Sales salaries 34900 34900 0 NONE
Depreciation 6600 6600 0 NONE
Insurance 1300 1300 0 NONE
Total fixed costs $42800 $42800 0 NONE
Total costs (33060+42800)= $75860 (34180+42800)= 76980 (75860-76980)= 1120 U

b)

Fallon Company
Selling Expense Flexible Budget Report
For the Month Ended March 31, 2020
Budget Actual Difference
Sales $181900 $181900 0 NONE
Less: Variable costs:
Sales commissions (181900*6%)= 10914 11340 (10914-11340)= 426 U
Advertising (181900*7%)= 12733 12280 (12733-12280)= 453 F
Travel (181900*4%)= 7276 6960 (7276-6960)= 316 F
Delivery (181900*2%)= 3638 3600 (3638-3600)= 38 F
Total variable costs $34561 $34180 (34561-34180)= 381 F
Less: Fixed costs
Sales salaries 34900 34900 0 NONE
Depreciation 6600 6600 0 NONE
Insurance 1300 1300 0 NONE
Total fixed costs $42800 $42800 0 NONE
Total costs (34561+42800)= $77361 (34180+42800)= $76980 (77361-76980)= 381 F